# Optimal DLT-based Solutions for the Perspective-n-Point

### v1

@@ I INTRODUCTION
This paper contributes to the state-of-the-art by improving the performance of the DLT for PnP, by leveraging ideas from optimal triangulation theory [22] and star-based attitude determination [23]. We show that the DLT can be weighted in a maximum likelihood sense if there is a priori information on the camera matrix. The a priori information can easily be obtained by a smaller-sized normalized DLT, RANSAC [24], or other inexpensive methods. Furthermore, we leverage the covariance information of the solution to weigh the orthogonal Procrustes problem [25] and retrieve an optimal $SO(3)$ rotation. While this two-shot approach comes at little extra computational cost for large $n$, it provides a dramatic improvement to the normalized DLT. Indeed, we find that our optimized DLT (oDLT) provides pose performance similar to the iterative methods but at the computational expense of the DLT.
@@ II STANDARD NORMALIZED DLT
where $\textbf{\emph{A}}\in\mathbb{R}^{2n\times 12}$. The least-squares solution to this homogeneous system is the null space of A, which corresponds to the one-dimensional subspace of the smallest singular value of A. Thus it may be found in the last column vector of V from the singular value decomposition (SVD)
@@ III OPTIMAL NORMALIZED DLT
$2\sum_{i}\text{vec}(\textbf{\emph{P}})^{T}\textbf{\emph{A}}_{i}^{T}\boldsymbol{\Sigma}_{\boldsymbol{\epsilon}_{i}}^{\dagger}\textbf{\emph{A}}_{i}=\mathbf{0}_{12\times 1},$ (13)
@@
which we rearrange to give,
@@
The system may be optimally solved accounting for uncertainties in both the pixel measurements ($\boldsymbol{\Sigma}_{\bar{\textbf{\emph{u}}}_{i}}$) and 3D points ($\boldsymbol{\Sigma}_{\bar{\textbf{\emph{p}}}_{i}}$), but for simplicity we proceed here by considering the pixel noise only. For isotropic noise ($\boldsymbol{\Sigma}_{\bar{\textbf{\emph{u}}}_{i}}=\sigma_{u}\textbf{\emph{S}}^{T}\textbf{\emph{S}}$), the pseudoinverse is analytically obtainable
@@ III-A Maintaining Optimality in $SE(3)$
where $\odot$ is the element-wise multiplication. This problem may be solved by converting the matrix representation to Rodrigues parameters and solving a weighted nonlinear least-squares ensuring $SO(3)$. We found that one iteration is enough for speed and accuracy.
r can be found back by applying the scale that ensures $det(\textbf{\emph{R}})=1$. However, its elements are not transformed in a way that maintains optimality. If the best position is needed, since the rotation has been recovered optimally, one can use LOST [22] as a non-iterative, DLT-based, and optimal $n$-view triangulation algorithm which minimizes reprojection error while maintaining $\mathcal{O}(n)$ complexity.
@@ III-C Summary: The Proposed Implementation
We propose the optimal DLT, oDLT in Algorithm. 1.
@@
3:Solve for an estimate of $\mathbf{P}$ using $\mathbf{A}$
@@
7:For better position, re-triangulate using LOST [22] (oDLT+LOST)
The algorithm may be written without any for loops in languages supporting linear algebra arithmetic.
@@ IV-A Numerical Experiments
In this numerical experiment, we compare the behavior of the DLT against different algorithms. For the DLT-based methods, we compare the regular un-normalized DLT DLT, the normalized DLT nDLT, the optimally weighted DLT oDLT, the optimally weighted DLT with a triangulation refinement oDLT+LOST, and the normalized DLT with a Gauss-Newton iteration from Ref. [14] nDLT+GN. We compare the DLT-based methods against some of the most popular methods — all of which are sourced from the publicly released code of the different references—, EPnP [10], EPnP+GN [10], CPnP [14], RPnP [12], OPnP [11]. The Gauss-Newton in EPnP+GN differs from a traditional Gauss-Newton as it only iterates on four parameters, and hence it is highlighted separately.
@@
Figure 3: Mean runtimes of all the algorithms.
RMSE $f_{x}$ RMSE $f_{y}$ RMSE $c_{x}$ RMSE $c_{y}$
nDLT 4.088 4.225 4.014 3.695
oDLT 3.917 4.060 3.823 3.516
TABLE I: Intrinsic parameter errors for $n$=50 (centered case).

### v2

@@ I INTRODUCTION
This paper contributes to the state-of-the-art by improving the performance of the DLT for PnP, by leveraging ideas from optimal triangulation theory [22] and star-based attitude determination [23]. We show that the DLT can be weighted in a maximum likelihood sense if there is a priori information on the camera matrix. The a priori information can easily be obtained by a smaller-sized normalized DLT, RANSAC [24] — with the benefit of providing inliers—, or other inexpensive methods. Furthermore, we leverage the covariance information of the solution to weigh the orthogonal Procrustes problem [25] and retrieve an optimal $SO(3)$ rotation. While this two-shot approach comes at little extra computational cost for large $n$, it provides a dramatic improvement to the normalized DLT. Indeed, we find that our optimized DLT (oDLT) provides pose performance similar to the iterative methods but at the computational expense of the DLT.
@@ II STANDARD NORMALIZED DLT
where $\textbf{\emph{A}}\in\mathbb{R}^{2n\times 12}$.
The least-squares solution to this homogeneous system is the null space of A, which corresponds to the one-dimensional subspace of the smallest singular value of A. Thus it may be found in the last column vector of V from the singular value decomposition (SVD)
@@ III OPTIMAL NORMALIZED DLT
$2\sum_{i}\text{vec}(\textbf{\emph{P}})^{T}\textbf{\emph{A}}_{i}^{T}\boldsymbol{\Sigma}_{\boldsymbol{\epsilon}_{i}}^{\dagger}\textbf{\emph{A}}_{i}=\mathbf{0}_{1\times 12},$ (13)
@@
which by using $\left[\textbf{\emph{P}}\bar{\textbf{\emph{p}}}_{i}\times\right]=\textbf{\emph{k}}^{T}\textbf{\emph{P}}\bar{\textbf{\emph{p}}}_{i}\left[\bar{\textbf{\emph{u}}}_{i}\times\right]$ (Eq. 1), yields
@@
This expression in Eq. 17 is very similar to the one obtained for triangulation in Ref. [28] for calibrated measurements. The system may thus be optimally solved accounting for uncertainties in both the pixel measurements ($\boldsymbol{\Sigma}_{\bar{\textbf{\emph{u}}}_{i}}$) and 3D points ($\boldsymbol{\Sigma}_{\bar{\textbf{\emph{p}}}_{i}}$), but for simplicity we proceed here by considering the pixel noise only. For isotropic noise ($\boldsymbol{\Sigma}_{\bar{\textbf{\emph{u}}}_{i}}=\sigma_{u}\textbf{\emph{S}}^{T}\textbf{\emph{S}}$), the pseudoinverse is analytically obtainable
@@ III-A Maintaining Optimality in $SE(3)$
where $\odot$ is the element-wise multiplication. We solve Eq. 29 by estimating a small rotation vector deviation $\delta\boldsymbol{\phi}$ from the nearest $SO(3)$ matrix to $\acute{\textbf{\emph{R}}}$, such that it minimizes this cost function. This may be done in a linear system by recognizing the small angle perturbation as $(\textbf{\emph{I}}-[\delta\boldsymbol{\phi}\times])$. Alternatively, this problem may be solved by using Rodrigues parameters in an iterative nonlinear least-squares ensuring $SO(3)$, where one iteration is typically enough.
r can be found back by applying the scale that ensures $det(\textbf{\emph{R}})=1$. However, its elements are not transformed in a way that maintains optimality. If the best position is needed, since the rotation has been recovered optimally, one can use LOST [22] as a non-iterative, DLT-based, and optimal $n$-view triangulation algorithm which minimizes reprojection error while maintaining $\mathcal{O}(n)$ complexity. Realizing that the weights in Eq. 24 are proportional to those in Ref. [22], we slice the system matrix $\textbf{\emph{A}}=[\textbf{\emph{B}},\textbf{\emph{C}}]$ to reformulate Eq. 25 as a triangulation problem,
$\begin{bmatrix}q_{1}\textbf{\emph{S}}\textbf{\emph{C}}_{1}\\ q_{2}\textbf{\emph{S}}\textbf{\emph{C}}_{2}\\ \vdots\\ q_{n}\textbf{\emph{S}}\textbf{\emph{C}}_{n}\end{bmatrix}\textbf{\emph{t}}=\begin{bmatrix}q_{1}\textbf{\emph{S}}\textbf{\emph{B}}_{1}\\ q_{2}\textbf{\emph{S}}\textbf{\emph{B}}_{2}\\ \vdots\\ q_{n}\textbf{\emph{S}}\textbf{\emph{B}}_{n}\end{bmatrix}\text{vec}(\textbf{\emph{R}}).$ (30)
Thus, given that we already have the optimal rotation R, we can easily estimate the optimal translation t and without recomputing the whole DLT system.
@@ III-B Comparing Our Approach to LM/GN
Different strategies could be used for the initial estimate. First, one could use a subset of the regular DLT system to solve for P, or use smaller-sized PnP methods such as P3P and P4P. Interstingly, if outlier rejection is a concern, the initial estimate could also come from a RANSAC scheme [24] along with the inlier indices and be seamlessly integrated with oDLT.
@@ III-C Summary: The Proposed Implementation
We propose the optimal DLT, oDLT in Algorithm. 1. The algorithm may be written without any for loops in languages supporting linear algebra arithmetic.
@@
3:Solve for an estimate of $\mathbf{P}$ using a subset of $\mathbf{A}$ (alternatively another fast PnP method, or RANSAC [24])
@@
7:For better position, re-triangulate using LOST in Eq. 30 (oDLT+LOST)
@@ IV-A Numerical Experiments
We compare the behavior of the DLT against different algorithms. For the DLT-based methods, we compare the regular un-normalized DLT DLT, the normalized DLT nDLT, the optimally weighted DLT oDLT, the optimally weighted DLT with a triangulation refinement oDLT+LOST, and the normalized DLT with a Gauss-Newton iteration from Ref. [14] nDLT+GN. We compare the DLT-based methods against some of the most popular methods — all of which are sourced from the publicly released code of the different references—, EPnP [10], EPnP+GN [10], CPnP [14], RPnP [12], OPnP [11]. The Gauss-Newton in EPnP+GN differs from a traditional Gauss-Newton as it only iterates on four parameters, and hence it is highlighted separately.
Figure 3: Mean runtimes of all the algorithms.
RMSE $f_{x}$ RMSE $f_{y}$ RMSE $c_{x}$ RMSE $c_{y}$
nDLT 4.088 4.225 4.014 3.695
oDLT 3.917 4.060 3.823 3.516
TABLE I: Intrinsic parameter errors for $n$=50 (centered case).
We only consider 2D noise and assume no covariance on the 3D points for simplicity of the framework. This assumption is consistent with all the other PnP methods listed in the prior paragraph. For full uncertainty awareness, Eq. 17 may be used with our method, in which case the results should be compared with MLPnP [18] and EPnPU [19].
@@ ACKNOWLEDGMENT
The authors would like to thank Lindsey Marinello and Tara Mina for their constructive feedback on this manuscript.